OpenStudy (anonymous):

What is the value for ƒ(x) = 4^2^x - 100 when x = 2 ?

OpenStudy (anonymous):

ƒ(x) = 4^2x - 100 when x = 2 f(x) = 4^2(2) - 100 f(x) = 16(2) - 100 f(x) = 32 - 100 f(x) = -68

OpenStudy (johnweldon1993):

\[\large f(x) = 4^{2^x} - 100 = (4^2)^x - 100= 4^{2x} - 100\] \[\large f(2) = 4^{2(2)} - 100\] \[\large f(2) = 4^{4} - 100\] \[\large f(2) = 256 - 100\] \[\large f(2) = 156\]